OpenStudy (anonymous):

Radical__?___ divided by 5 = 2. What goes in the radical?

terenzreignz (terenzreignz):

It looks like I'm still here after all... It's like this, see \[\Large \frac{\sqrt x}{5}=2\]

OpenStudy (anonymous):

Yes thats what it looks like but I dont know how to solve for x though

OpenStudy (anonymous):

And thank you for helping me

terenzreignz (terenzreignz):

Yeah, well, that's why I'm here :) multiplying both sides by 5 will cancel the denominator on the left, right?

OpenStudy (anonymous):

Im not sure.. I dont know how to do this :/

terenzreignz (terenzreignz):

Okay, keep calm and carry on ;) \[\Large 5 \times \frac{\sqrt x}{5}= 2 \times 5\] right?

OpenStudy (anonymous):

There has to be an answer for x that will = 2

terenzreignz (terenzreignz):

Yeah, cool it :) I just multiplied both sides by 5... basic Multiplication property of equality; if two sides are equal, multiplying both sides of the equation by the same number would not affect the equality...

OpenStudy (anonymous):

Okay thank you :D

terenzreignz (terenzreignz):

So, 5 cancels out at the left... \[\large \cancel 5 \times \frac{\sqrt x}{\cancel 5}= 10\] \[\Large \sqrt x = 10\] can you do it from here?

OpenStudy (anonymous):

Yes thank you

terenzreignz (terenzreignz):

Okay, good :) What's your answer?

OpenStudy (anonymous):

Hold on still trying to figure it out

terenzreignz (terenzreignz):

Squaring cancels out the square root... in that... \[\Large \sqrt{a}^2 = a\]

terenzreignz (terenzreignz):

So we can square both sides of the equation... \[\Large \sqrt x^2 = 10^2\] we get \[\Large x = 100\] And that's it :) Don't worry, in time, this will be second nature to you :)

OpenStudy (anonymous):

oh my goodness! i dont understand math

terenzreignz (terenzreignz):

But you understand how I got the answer? :)

OpenStudy (anonymous):

Ys I do Ill try using that for when another question pops up like that

terenzreignz (terenzreignz):

Sorry, it might have been too quick... a bit crash-course - y

terenzreignz (terenzreignz):

I have to go now, okay? Have fun with Math ^.^ ----------------------------------------------- Terence out

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

I understand what you werre trying to show me :) I tried a couple of problems and I finaly got it down :D thanks soooo much!